OpenStudy (anonymous):

Rewrite the following parametric equations by eliminating the parameter. x = t y = -t2 + 3 Would my answer be y=-x^2+3 or y=x2+3

OpenStudy (anonymous):

sorry I mean y=-x^2+3 or y=x^2+3